### v1

# Exploring run-and-tumble movement in confined settings through simulation

Abstract
Motion in bounded domains is a fundamental concept in various fields, including billiard dynamics and random walks on finite lattices, with important applications in physics, ecology and biology. An important universal property related to the average return time to the boundary, the Mean Path Length Theorem (MPLT), has been proposed theoretically and confirmed experimentally in various contexts. In this discussion, we investigate a wide range of mechanisms that lead to deviations from this universal behavior, such as boundary effects, reorientation and memory processes. In particular, this study investigates the dynamics of run-and-tumble particles within a confined two-dimensional circular domain. Through a combination of theoretical approaches and numerical simulations, we validate the MPLT under uniform and isotropic particle inflow conditions. The research demonstrates that although the MPLT is generally applicable for different step length distributions, deviations occur for non-uniform angular distributions, non-elastic boundary conditions or memory processes. These results underline the crucial influence of boundary interactions and angular dynamics on the behaviour of particles in confined spaces. Our results provide new insights into the geometry and dynamics of motion in confined spaces and contribute to a better understanding of a broad spectrum of phenomena ranging from the motion of bacteria to neutron transport. This type of analysis is crucial in situations where inhomogeneity occurs, such as multiple real-world scenarios within a limited domain. This bridges the gap between theoretical models and practical applications in biological and physical systems since the study of the statistics of movement in confined settings can bring some light in explaining the mobility mechanism of active agents.
1 Introduction
The properties of motion within bounded domains are applicable across a variety of fields such as physics, ecology and biology. Motion in bounded domains has far-reaching implications, from fundamental mathematical insights to practical applications in environmental science, biology, medicine and ecology. It lays the foundation for predicting and controlling the behavior of particles or agents in diverse and restricted environments, offering potential benefits to numerous scientific and technological fields.
A key aspect of motion in confined environments is the Mean Path Length Theorem (MPLT), a generalization of a geometric result originally proposed by Cauchy [1]. The (MPLT) is a fundamental result in the study of particle dynamics in confined domains, providing a universal expression for the average path length of particles. In particular, this theorem states that the mean path length $<L>$ of particles within a bounded domain $\Omega$ is determined by the ratio between the volume $V_{\Omega}$ and the surface area $\Sigma_{\Omega}$, regardless of the details of the dynamics of the particles or the shape of the domain.
$<L>=\eta_{d}\frac{V_{\Omega}}{\Sigma_{\Omega}},$ (1)
where $\eta_{d}$ is a numerical factor depending only on space dimensionality $d$ (in the case we will consider in detail $d=2$ and $\eta_{2}=\pi$).
More precisely, the MPLT concerns the case where $<L>$ is the mean length of randomly distributed chords intersecting a convex body. Based on the linear Boltzmann equation, the MPLT was also extended to scenarios in which the particles do not move in straight lines but are scattered [2]. This showed that the Cauchy formula applies more generally to the stochastic paths of Pearson random walks with exponentially distributed flight lengths. This demonstrates that under the conditions of a uniform and isotropic particle influx, the absence of absorption or particle reproduction within the medium, and a detailed balance in scattering probabilities, the average path length remains unchanged. The analysis contributes to the understanding of transport phenomena in scattering media, which are relevant, for example, for the estimation of reaction rates in neutron transport [3, 4, 5], ionic recombination in gasses and other processes in finite volumes. Furthermore, this theorem has recently been generalized for non-convex domains and extended to the case of Brownian motion [6, 7, 8]. The work of Benichou et al. [9] discusses the averaged residence times of stochastic motions within bounded domains, focusing on the mean first exit and residence times for particles undergoing random motions, such as Pearson random walks, by employing a backwards Chapman Kolmogorov equation. This study is motivated by various physical systems, including animal trajectories [10] and neutron scattering processes. In [11] the case of arbitrary closed trajectories in arbitrary domains is investigated and it is shown that the mean path length still agrees with the Cauchy formula as long as no closed trajectory is completely contained in the domain. Shukla et al. [12] extend the Cauchy formula to lattice-based random walks, providing a straightforward understanding compared to previous models relying on continuum-based Boltzmann transport equations. By analyzing random walks on an $N\times N$ square lattice, the research provides both numerical results and theoretical underpinning of the observed phenomena, suggesting that the Cauchy formula can be generalized beyond traditional applications to include more complex random walk scenarios. This property holds for anomalous and classical photon propagation [13]. All these generalizations imply that the Cauchy formula, Eq. (1) is applicable to a wide range of scenarios and phenomena, hence the importance. Cauchy’s universality was recently confirmed by experimental results on the movement of bacteria [14] and the propagation of light in scattering media [15, 16].
In addition, to know the statistics of random walks in confined settings offer insights into a range of phenomena, such as the conditions under which active particles can escape barriers or be trapped by external forces, or how chemotactic fields influence particle motion towards or away from high nutrient concentrations, applicable to bacterial motion, intracellular transport, and animal foraging, bridging the gap between theoretical models and real-world biological processes [17, 18, 19]. These systems presents what is called a Run-and-tumble movement, characterized by randomly reoriented ballistic movements. Understanding the mean chord length can provide foundational insights into the geometry and scale of motion within a domain, which is indirectly relevant to estimating escape times or paths in narrow escape scenarios, especially when considering the impact of domain geometry on escape dynamics [20, 21, 22] or target searching [23].
It is widely recognized that numerous stochastic transport phenomena observed in the real world do not exhibit jump lengths that are exponentially distributed, nor do they show scattering angles that are uniformly random [25, 26, 27]. For instance, isotropic random flights having fixed length have application in polymer chains [28] and an exponential law has application, as mentioned, to neutron diffusion processes and locomotion problems in biology. It naturally raises the question of whether these processes still adhere to similar universal characteristics. In fact, heuristic reasoning indicates that Cauchy’s formula is applicable to any (isotropic) random walks as long as the random walk’s entry into the domain involves a length distribution that aligns with equilibrium [9, 29, 30]. In particular, some attention was recently given in the literature to the role of boundary conditions in influencing the absorption time of run-and-tumble particles [31, 32], and to the particle’s speed distribution or the spatial dimensionality [33].
Complex dynamics in bounded domains are occasionally linked to non-homogeneous distributions, as in the case of active particles (see, for example, [34, 35, 36]) or when interactions with walls are present [37]. Therefore, it is natural to wonder if Cauchy universality is upheld in these situations. This is the main topic we will cover, giving examples of mechanisms causing nonuniform stationary probability distributions and breaking Cauchy universality; we will take into account both the scenario where particles move in a stochastic way inside the body (random walk) and the scenario where they follow rectilinear trajectories between consecutive collisions with the boundary (billiard). Through theoretical analysis and numerical simulations, this study examines how different boundary and reorientation conditions, affect the motion’s universal properties. This research offers potential implications for understanding motion in confined systems across various scientific domains, extending beyond the classical applications of billiards and random walks to include biological and physical phenomena where such conditions are prevalent.
The paper is structured as follows. In the next section, the methodology for simulating particle movement within a 2D circular space is described, including variations in step length and angular distributions under different boundary conditions. Section 3 presents the analysis confirming the MPLT under uniformly random angular distributions and highlights deviations with non-uniform distributions. The section 4 examines path length distributions and their variance, illustrating the transition from diffusive to quasi-ballistic movements. Section 5 explores the radial probability distributions, showing uniformity under elastic boundary conditions and deviations with random angle conditions. In section 6, we introduce memory effects in particle reorientation and their impact on path length and radial distributions. Finally, we summarize the key findings and their implications in section 7.
2 Simulations
Figure 1: Trajectory of the particle inside a circle with elastic boundary. Different types of movement can be obtained by changing the ratio between the mean step length and the circle radius, $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. Figure (a) is diffusive, on the other extreme, (d) is quasi-ballistic.
To replicate a simple geometric setting common to experimental realizations [18, 19, 14, 37], our code simulates a particle moving within a unit 2D circular space. In our model, the particle travels at a constant unit speed, $v=1$. Therefore, the time between successive collisions is equal to the total trajectory length, allowing us to discuss average residence time and mean path length interchangeably. In two dimensions, the mean path length, as defined in eq. (1), is expressed as follows:
$<L>=\pi\frac{S}{P},$ (2)
where S is the surface and P the perimeter of the domain. So, for the disk of radius $R=1$ we have
$<L>=\frac{\pi}{2}R=\frac{\pi}{2}.$ (3)
After each step (run), the particle changes direction (tumble). Consequently, to define the Run-and-Tumble dynamics, we must specify: (i) the distribution of step lengths, $Q(\ell)$; (ii) the angular distribution of the tumble events; and (iii) the outgoing angle after a collision with the boundary. In this work, we explore the statistics of the run-and-tumble movement under four conditions of step length distribution, namely: (a) Gaussian distributed (as in some standard random walks),
$Q(\ell)=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{1}{2}(\frac{\ell-l}{\sigma})^{2}\right),$ (4)
where $l=<\ell>$ is the mean step length and $\sigma$ is the standard deviation, (b) uniformly random, (c) fixed (as in polymer chains [28]), and (d) exponential distributed (typical of bacteria mobility [14] and neutron scattering [3]),
$Q(\ell)=\frac{1}{l}\exp\left(-\frac{\ell}{l}\right).$ (5)
In the case of the Gaussian distribution, no cutoff was used. This means that there is a possibility that the particles move backwards if $\ell$ takes a negative value. We will analyze the statistics and the validation of the MPLT under various angular distributions: (a) a fixed angle, which can be considered as a representation of strong chemotaxis (preferred direction); (b) a Gaussian distribution, representing weak chemotaxis, since there is a preferred direction with some deviation, a similar approach is used in [38]; (c) the particle is allowed to move in any direction uniformly distributed in $[0,2\pi]$, which is the typical case considered in most studies.
When the particle reaches the boundary, it changes its direction of motion as a result of the collision, and the new direction may depend on the angle of incidence. In this paper, we will consider two different boundary conditions: 1- elastic (i.e., the angle of reflection is equal to the angle of incidence), 2- the particle exits the collision with the boundary at a random outgoing angle.
By adjusting the mean step length, $l$, we can observe different types of movement, as illustrated in Fig. (1). While these examples start from the center of the circle, the findings presented in this work are not dependent on the starting location, including points on the boundary. In Fig. (1), the exit angle after each tumble event is uniformly random, and the step length follows an exponential distribution. We observe various movement types according to the ratio $l/R$, where $R$ is the circle’s radius. When $l/R<<1$, the movement is diffusive, but when $l/R>>1$, the movement becomes quasi-ballistic. Purely ballistic movement occurs in the limit as $l/R\to\infty$.
3 Mean path length theorem
First, we explore the conditions under which the Mean Path Length Theorem (MPLT) is valid. To do this, we calculate the mean path length using various combinations of step length and angular reorientation distribution. This analysis is conducted with elastic boundary conditions for four distinct values of $l/R$: $0.01$, $0.1$, $1$, and $10$. We find that the MPLT is valid only when the angular distribution is uniformly random, across all step length distributions. At first, the fact that the mean path length is the same for different values of $l/R$ seems very surprising, because if we replace straight segments (or arcs) by irregular paths, we intuitively expect a twofold mechanism to profoundly change the dynamics (stochastic short returns to the boundary and long walks in the domain without touching the boundary) [29]. This fact is confirmed by plotting the ratio $(<L>_{s}/<L>_{t})-1$, where $<L>_{s}$ represents the mean path length from the simulation, and $<L>_{t}$ is the theoretical value predicted by the MPLT. If the MPLT is accurate, the simulated and theoretical mean path lengths will converge as the number of collisions increases. Here, we present some of the most illustrative cases. In Fig. (2), a typical graph of $(<L>_{s}/<L>_{t})-1$ is shown for various $l/R$ ratios, using an exponential step length distribution and a uniformly random angular distribution. It is observed that the difference between simulated and theoretical values tends to zero as the number of collisions increases.
Figure 2: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic boundary condition, exponential step length distribution, and uniformly random angular distribution. Each point is an average of 20 simulations. It can be seen that they coincide when the number of collisions approach to infinity.
The step length distribution can be adjusted to any of the mentioned scenarios, and the MPLT will still hold true. For illustration, in Figs. (3(a)) and (3(b)), we present two distinct cases of step length distribution: (a) random between 00 and $2l$, and (b) a constant value of $l$. The differences between the simulated and theoretical values of mean path length approach zero as the number of collisions tends towards infinity in all instances.
Figure 3: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic conditions at the boundary, random angle distribution, and (a) random step length between 0 and $2l$, (b) constant step length $l$.
On the other hand, if we alter the distribution of the reorientation angle after the tumble events, the assumption of homogeneous flux at the boundary no longer applies, and consequently, the mean path length does not converge to the same value for all $l/R$ ratios, as illustrated in Figs. (4(a)) and (4(b)). In Fig. (4(a)), the angles after the tumble event follow a Gaussian distribution with a mean of 00 and a standard deviation of $\pi/2$. We observed similar outcomes with various combinations of means and standard deviations. Similarly, Fig. (4(b)) displays results for a fixed angle of $\pi/4$, yet comparable results are achieved with different preferred direction values.
Figure 4: Difference between predicted value of the mean path length according to the theorem and the value of the simulation versus the number of collisions with elastic conditions at the boundary, exponential step length distribution, and (a) gaussian angle distribution with mean 0, and $\sigma=\pi/2$; (b) fixed value of angle at $\pi/4$.
The fixed angle and the Gaussian distribution around an angle create a preferred direction, observable in some experimental behaviors of bacteria due to mechanisms like chemotaxis. The fixed angle constitutes a strong condition (the particle consistently moves in a specific direction), while the Gaussian distribution signifies a less stringent condition (there is a preferred direction, yet a certain probability exists for the particle to move in another direction close to the preferred one). In scenarios with fixed and Gaussian distributed angles, the particle tends to collide with the boundary consistently in the same region, as illustrated in Fig. (5). This results in non-uniformity and anisotropy, thus violating the assumptions underlying the MPLT. This result is not surprising, as previous works had already clarified the conditions of the angular distribution under which the theorem holds [2, 7, 9, 30].
Figure 5: Trajectory of 1000 steps of the particle inside a circle with elastic boundary, exponential step length distribution, and $l/R=0.1$. (a) The angle distribution is normal distributed around $\theta=0$, (b) Angle distribution fixed at $\pi/4$. The particle has a strong tendency to moves always to the same part of the border, producing anisotropy and therefore breaking the hypothesis of the MPLT.
Let’s now examine the scenario where the outgoing angle after a collision with the boundary is random. Initially, the absolute values of the mean path length under elastic boundary conditions, with an exponential step length distribution and uniformly random angle distribution, are presented in Fig. (6(a)). These values approach the theoretical value of $\frac{\pi}{2}=1.57079...$ as the number of collisions approaches infinity. For contrast, the absolute values of the mean path length for the random reflecting angle condition are plotted in Fig. (6(b)). As can be seen, the invariant property does not hold anymore, as we anticipated in [32]. On the contrary, the mean path length does depend on $l/R$, i.e., the type of movement (diffusive or ballistic), as can been appreciated in Fig. (6(b)).
Figure 6: (a) Mean path length versus the number of collisions with elastic reflecting conditions at the boundary. Each point is an average of 20 simulations. It can be seen that they approach to the theoretical value as the number of collisions tends to infinity. (b) Mean path length versus the number of collisions with random reflecting conditions at the boundary. Each point is an average of 20 simulations. In this case, the mean path length in not invariant and depends on the value of $l/R$.
In Fig. (7), we show this dependency in a $<L>$ versus $l/R$ plot for simulations with $10^{9}$ collisions. It can be seen that the mean path length tends to $4/\pi$, which is the mean path length of a random billiard as showed in [32].
Figure 7: Mean path length as a function of the parameter $l/R$ for simulations with $10^{9}$ collisions and random angle boundary conditions. The limit for $l/R\rightarrow\infty$ is the mean path length of a random billiard, $1.27324/\pi=$ 1.2732 ….
The MPLT holds true across various step length distributions -whether Gaussian, uniformly random, fixed, or exponentially distributed- as these do not affect the uniformity of particle incidence on the boundary, a critical assumption of the theorem. However, altering the angular distribution of the tumble events or the angle post-collision breaks this assumption. When the angular distribution deviates from being uniformly random, it introduces anisotropy or non-uniformity in the way particles interact with the boundary, violating the MPLT’s conditions. For instance, fixed or Gaussian-distributed angles after tumble events create preferred directions of movement, leading to non-homogeneous wall collisions and thereby the theorem’s failure. The MPLT does not hold under random boundary conditions, as opposed to elastic conditions, due to the way these conditions alter the uniformity and isotropy of particle interactions with the boundary of the domain. On the contrary, in elastic boundary conditions, a particle’s angle of reflection equals its angle of incidence, preserving the distribution of angles at which particles strike and leave the boundary. This uniformity ensures that the particle flux across the boundary maintains the assumptions required for the MPLT to apply.
4 Distributions of path length between collisions
It is evident that under the condition of homogeneous and isotropic flux of particles at the boundary, the mean path length is determined by Cauchy’s formula. However, it prompts the question: how does the distribution behaves in different conditions? Let us delve into this inquiry.
Figure 8: (a) Variance of the path length versus the number of collisions with elastic reflecting conditions at the boundary. (b) Variance of the path length versus the parameter $l/R$ for simulations with $10^{9}$ collisions with elastic and random reflecting conditions at the boundary in log-log representation.
Hereinafter, we will focus in the particular case of exponential step length distribution and uniformly random angular distribution. In the subsequent analysis, we will further explore the dynamic properties of the path length distributions by examining the variance of these distributions. The variance provides crucial insight into the spread of the path length around its mean, which in this context is constant across all distributions at $<L>=\frac{\pi}{2}$ for the elastic boundary conditions, and variable as showed in Fig. (7) for the random angle boundary condition. In Fig. (8), we plot the variance of the path length as a function of two parameters: the number of collisions for the elastic condition, and the parameter $l/R$ for both, elastic and random boundary conditions. This plot illustrates how the variance evolves as the motion transitions from diffusive to quasi-ballistic regimes across different scaling of $l/R$. The limit as $l/R\rightarrow\infty$ is $Var(L)_{elastic}=8/3-\pi^{2}/4\sim 0.199$ for the elastic case. On the contrary, the random billiard has the limit $Var(L)_{random}=2-16/\pi^{2}\sim 0.378$. These limits can be obtained by a straightforward calculation. For the elastic case, the measure is $\mu(d\theta)=\frac{1}{2}\cos\theta d\theta$, and, by the geometry of the problem, a chord between two points in the boundary has length $L=2R\cos\theta$, see Fig. (14(a)), where $\theta$ is the outgoing angle from the first point to the second one, measured with respect to the normal. Therefore,
$<L^{2}>_{elastic}=\int_{-\pi/2}^{\pi/2}(2R\cos\theta)^{2}\times\frac{1}{2}\cos\theta\times d\theta=\frac{8}{3}R^{2},$ (6)
and the variance for our case $R=1$ is $Var(L)_{elastic}=<L^{2}>-<L>^{2}=8/3-\pi^{2}/4$. On the other hand, in the case of random boundary condition, the measure is $\mu(d\theta)=d\theta/\pi$, corresponding to a collision rule $P_{R}(\theta)=1/\pi$, the first moment is $<L>_{random}=4/\pi$ [32], and the second moment is
[…]
Therefore the variance (with $R=1$) is $Var(L)_{random}=2-16/\pi^{2}$.
Figure 9: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. It can be seen how the distribution evolves from diffusive (a) to quasi-ballistic (d).
A more clear picture can be made by looking at the distributions. In Fig. (9), the distribution of path length between collisions is shown, featuring elastic boundaries, and the conditions previously mentioned, for various values of the parameter $l/R$: (a) $l/R=0.01$, (b) $l/R=0.1$, (c) $l/R=1$, (d) $l/R=10$. As illustrated in Fig. (2), while all distributions share the same mean, $<L>=\pi/2$, one can observe how the distribution transitions from diffusive (a) to quasi-ballistic (d).
Figure 10: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. Ratio $l/R=0.1$. (a) linear, (b) semilog, and (c) log-log representation. Red lines are exponential fittings, green lines are power law fittings. It can be appreciated that the distribution is a combination of both functions.
Interestingly, when $l/R<1$, the distribution is an exponential decay with a power law deviation for small values of length, as illustrated in Fig. (10) and (11). The distributions are presented in (a) linear, (b) semilog, and (c) log-log formats. It is noteworthy that for $l/R=0.3$, the semilog representation displays a straight line, whereas for $l/R=0.05$, it is the log-log representation that appears straight. If we decrease the ratio $l/R$, the power law deviation for small lengths grows.
Figure 11: Path length distribution with elastic boundaries, $10^{7}$ collisions, 100 bins. Ratio $l/R=0.3$. (a) linear, (b) semilog, and (c) log-log representation. The red line is an exponential fitting of the distribution.
Figure 12: Cumulative probability distributions of path length between collisions for (a) $l/R=0.002$ and (b) $l/R=1$. Both with elastic boundary conditions, exponential step length, uniformly random angular distribution, $10^{8}$ collisions, and 600 bins.
This is better appreciable by looking at the cumulative distributions. We show the cumulative probability distribution in Fig. (12) for two different ratios $l/R$. The distributions were made with $10^{8}$ collisions and 600 bins. It is evident that, excluding small lengths, an exponential decay, $P=P_{0}\exp(-\gamma\ell)$, can be fitted. The parameter of the decay, $\gamma$, depends on $l/R$, as shown in Fig.(13). A parabolic fit seems to represent the numerical data quite well, but we have no theoretical argument to support this choice.
Figure 13: Exponent $\gamma$ of the exponential fit of cumulative distributions as a function of $l/R$. The red line is a parabolic fit of the points.
When $l/R>1$, the distributions, Fig. (9(d)) for example, are a combination of a diffusive and a ballistic distribution, since sometimes the step length is equal or greater than $R$, contributing to the ballistic mode, and other times the step length is small and therefore contributes to the diffusive mode. Let us name $\tau$ the length of a chord between two points of the boundary. Remembering that $\tau=2R\cos\theta$, see Fig. (14(a)), where $\theta$ is the outgoing angle from the first point to the second one, measured with respect to the normal. Then we have
$|\frac{d\tau}{d\theta}|=2R\sin\theta=\sqrt{4R^{2}-\tau^{2}},\hskip 8.53581pt\theta\in[0,\pi/2].$ (8)
When the boundary conditions are elastic, the distribution verifies
$\Pi(\tau)d\tau=\cos\theta d\theta,$ (9)
$\Pi(\tau)=\frac{\cos\theta}{|d\tau/d\theta|}=\frac{\tau}{2R\sqrt{4R^{2}-\tau^{2}}}.$ (10)
As can be seen in Fig. (14(b)), the ballistic part can be descripted with high accuracy by eq. (10). In this figure, we show the ballistic part of the path length distribution for elastic boundary conditions, exponential step length distribution and uniformly random angular distribution. The distribution is constructed with 100 bins and $10^{7}$ collisions, for the case $l/R=10$.
Figure 14: (a) Geometry of a chord. (b) Ballistic part of the path length distribution with elastic boundaries. In black: simulation for $l/R=10$, $10^{7}$ collisions, and 100 bins. In red: theoretical model, eq. (10).
5 Radial distributions
Let us now study the radial distributions, i.e., the probability of finding a particle at a given distance r from the center of the confined domain, $p(r)$. This distribution helps in understanding how particles spread out within a bounded area, which is essential for predicting escape times, collision frequencies, and interaction rates with boundaries. The radial distribution of run-and-tumble particles in a confined two-dimensional domain is a critical aspect that provides insight into the spatial dynamics and density distribution of particles within the domain. This section explores the relevance of radial distribution in understanding run-and-tumble movement.
We performed the simulations with exponential step distribution and uniformly random angular distribution. The radial distribution was found to be linear under elastic boundary conditions for all values of the ratiof $l/R$. For instance, in Fig. (15(a)), we show the distribution for the case $l/R=0.3$. This can be analytically calculated as follows. Imagine a chord between points at the boundary representing a part of the trayectory of the particle, as represented in Fig. (14(a)). The probability of being at a given point on the chord, $z$, at a distance $r$ from the center is
$W_{\theta}(r)dr=\frac{dz(r)}{2R\cos\theta},$ (11)
where $\theta$ is the angle measured with respect to the normal, see Fig. (14(a)). By geometry we get
$r^{2}=z^{2}+R^{2}\sin^{2}\theta,$ (12)
and therefore,
$\frac{dz}{dr}=\frac{2r}{\sqrt{r^{2}-R^{2}\sin^{2}\theta}},$ (13)
$W_{\theta}(r)=\frac{1}{R\cos\theta}\frac{r}{\sqrt{r^{2}-R^{2}\sin^{2}\theta}}.$ (14)
Taking into account that the measure is $1/2.\cos\theta.d\theta$ and the chord length is $\tau=2R\cos\theta$, we get that
$<\tau>=R\int_{-\pi/2}^{\pi/2}cos^{2}\theta.d\theta=\frac{\pi R}{2},$ (15)
and therefore, the radial probability can be calculated as
$p_{elastic}(r)=\int\frac{\tau}{<\tau>}W_{\theta}(r)\cos\theta.d\theta=\frac{4r}{\pi R}\int_{0}^{arcsin(r/R)}\frac{\cos\theta.d\theta}{\sqrt{r^{2}-R^{2}sin^{2}\theta}}$ (16)
Making a change of variables,
$\sin\phi=\frac{R}{r}\sin\theta$ (17)
$\cos\phi.d\phi=\frac{R}{r}\cos\theta.d\theta$ (18)
We obtain,
$p_{elastic}(r)=\frac{4r}{\pi R}\int_{0}^{\pi/2}\frac{\cos\phi.d\phi}{r\cos\phi}=\frac{2r}{R^{2}}.$ (19)
The last equation implies a uniform invariant density, $\rho$, since
$\rho_{elastic}=\frac{p_{elastic}(r)}{2\pi r}=\frac{1}{\pi R^{2}}.$ (20)
This theoretical model matches well with the simulation results, confirming the uniform distribution of particles when the boundary conditions are elastic.
Figure 15: (a) Radial distribution elastic boundary conditions, exponential step length and uniformly random angular distribution. In black: simulation for $l/R=0.3$, $10^{7}$ collisions, and 100 bins. In red: theoretical model, eq. (19). The linear dependence indicates that there is a uniform invariante density. (b) Radial probability distribution for random reflecting angle at the boundary. In black: simulation for $l/R=10$, $10^{7}$ collisions, 100 bins. In red: theoretical model, eq. (23).
In contrast, for random angle boundary conditions, the radial distribution deviates from linearity and is described by the complete elliptic integral of the first kind that can be calculated as follows. In the case of random angle boundary conditions, the measure is $d\theta/\pi$, and therefore,
$<\tau>=\frac{2R}{\pi}\int_{-\pi/2}^{\pi/2}cos\theta.d\theta=\frac{4R}{\pi}.$ (21)
The radial distribution is
$p_{random}(r)=\int\frac{\tau}{<\tau>}W_{\theta}(r)\frac{2}{\pi}.d\theta=\frac{r}{R}\int_{0}^{arcsin(r/R)}\frac{d\theta}{\sqrt{r^{2}-R^{2}\sin^{2}\theta}}.$ (22)
Using the same change of variable as before we get,
$p_{random}(r)=\frac{r}{R^{2}}\int_{0}^{\pi/2}\frac{d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=\frac{r}{R^{2}}K\left(\frac{r}{R}\right),$ (23)
where $K(x)$ is the complete elliptic integral of the first kind, defined as [39]
$K(x)=\int_{0}^{\pi/2}\frac{ds}{\sqrt{1-x^{2}sin^{2}s}}.$ (24)
This expression is compared with the simulation in Fig. (15(b)) for the case $l/R=10$ and, as can be seen, the agreement is satisfactory. An alternative procedure for calculating these distributions are presented in the appendix. This distribution shows that particles are more likely to be found near the walls of the domain under random angle reflections, these results seem to be in agreement with experiments with microbia in which the microbe spends the majority of their time close to the walls of the domain [37, 40]. In fact, the study of the radial distribution can bring information about the mobility mechanism in experimental settings, since different distributions are generated depending on how the active agent interacts with the walls and, as we will show in the next section, the reorientation mechanism. It would be interesting, therefore, to see more experimental outcomes on this matter. A related concept, the boundary local time, i.e. how long a diffusing particle spends in a close vicinity of a confining boundary, plays thus a crucial role in the description of various surface-mediated phenomena, such as heterogeneous catalysis, permeation through semipermeable membranes, or surface relaxation in nuclear magnetic resonance [41].
6 Processes with memory
Let us now focus in a different, non-trivial case. Process memory refers to the phenomenon where the future state of a particle’s motion is influenced by its past states, leading to non-Markovian dynamics. The concept of process memory plays a crucial role in understanding the reorientation mechanism of active agents in run-and-tumble movement. For example, some processes like chemotaxis, introduce a kind of memory, namely, the ability of the particle to compare the concentration or the gradient of a chemical in different steps of its movements.
Figure 16: Mean path length (a) and variance of path length (b) versus $l/R$ for three values of constant drift in the reorientation angle for simulations with $10^{9}$ collisions, elastic boundary, and exponential step length distribution.
Process memory can significantly alter the motion patterns of particles, leading to deviations from theoretical predictions like the MPLT. We are going to introduce a simple memory process in the simulations, following the next rule: the new reorientation angle of the tumble events is equal to the last one plus a constant, $\theta_{n}=\theta_{n-1}+C$. The step length distribution is exponential and the border elastic. We explore the mean and variance of the distributions for three different values of the constant $C$, as shown is the Fig. (16).
Figure 17: Path length distribution for a process with memory, elastic boundaries, exponential step length distribution, $10^{7}$ collisions, 100 bins. Constant drift $C=\pi/8$. (a) $l/R=0.3$, (b) $l/R=1$, and (c) $l/R=5$.
The limit value of the mean path length at $l/R\rightarrow\infty$ is the Cauchy prediction, $\pi/2$, for all values of constant $C$. This occurs because the movement is essentially a billiard, and the influence of the memory process vanishes. The variance decreases because the probability distribution narrows by $l=2R$, as can be seen in Fig. (17). In this type of memory process, the distribution has two peaks, one of which decreases and the other increases as $l/R$ increases. This shows that memory processes introduce different movement patterns that deviate from the expected behavior under Markovian assumptions.
Figure 18: Radial distribution for a process with memory, elastic boundaries, exponential step length distribution, $10^{7}$ collisions, 100 bins. Constant drift $C=\pi/8$. (a) $l/R=0.1$, (b) $l/R=0.5$, and (c) $l/R=1$.
Regarding the radial distribution, this presents a non-trivial form, as appreciated in Fig. (18). This shows how the reorientation mechanism of an active particle can produce distinctive distributions, hence bringing information about its underlaying laws.
Understanding process memory is not only theoretically intriguing but also practically significant. For instance, [14] demonstrated invariance properties of bacterial random walks in complex structures, suggesting that real-world bacterial motion often exhibits memory effects that need to be accounted for in predictive models. Additionally, the impact of memory on narrow escape times and sorting of active particles was explored in [20], emphasizing the practical applications in microfluidic device design and biological systems .
By integrating these insights, it becomes clear that memory effects play a fundamental role in the dynamics of confined run-and-tumble particles. This study’s exploration of simple memory processes provides a foundational understanding that can be expanded with more complex, real-world inspired mechanisms in future research. Such efforts will enhance our ability to predict and control particle behavior in various scientific and technological applications.
7 Conclusions
In this paper, we explored the dynamics of run-and-tumble motion within confined two-dimensional domains, specifically examining how various distributions and boundary conditions affect the validity of the Mean Path Length Theorem (MPLT) and the statistics of the dynamics. Through a combination of theoretical analysis and numerical simulations, we demonstrated that the MPLT holds true across a wide range of conditions, provided that the angular distribution of particle tumbling remains uniformly random. This establishes the significance of understanding motion within bounded domains across physics, ecology, and biology, and highlights the foundational role of the Mean Path Length Theorem (MPLT) in predicting and controlling particle or agent behavior in diverse environments.
We used a simulation of a 2D circular domain to model particle movement, adhering to real-world experimental conditions. By employing a constant speed for particles and exploring various step length distributions, we observe the transition of movement from diffusive to quasi-ballistic. Then, we presented an analysis on the MPLT’s applicability under diverse conditions, revealing that the theorem holds true when angular distributions are uniformly random. In fact, when particles exhibit a strong directional bias due to fixed or Gaussian-distributed angles, they tend to collide with the same region of the boundary, introducing anisotropy and thereby violating the MPLT’s foundational assumptions.
We examined path length distributions and their variance under different conditions. Exploring path length distributions between collisions sheds light on the transition from diffusive to quasi-ballistic movements. This part of the study illustrates how varying the parameter $l/R$ influences the distribution, emphasizing the intricate relationship between movement type and path length distribution. The results show a transition from diffusive to quasi-ballistic regimes as $l/R$ increases. For $l/R<1$, the distribution is characterized by an exponential decay with a power-law deviation for small path lengths, transitioning to a combination of diffusive and ballistic behaviors as $l/R$ increases. Understanding these distributions helps in predicting escape times, optimizing sorting mechanisms, and designing environments that mimic natural habitats for microorganisms.
The analysis of radial distributions under different boundary conditions reveals a uniform density across all cases of $l/R$ when conditions are elastic, demonstrating a consistent linear radial distribution. The contrast with random angle boundary conditions, where the radial distribution deviates, highlights the significant impact of boundary interactions on particle distribution. This can be relevant to other physically important problems of motion in bounded systems, such as the problem of narrow escape [42], since increased spatial density near the edge increases the probability of hitting the target. Random reflections can also be important when considering reversible sticking to the boundary (see e.g. [43, 44, 45]), because if a particle sticks to the boundary, it is natural to consider the outgoing angle of release as random and uncorrelated with the incoming direction. We show in this work that the study of the statistics of movement in confined settings can bring some light to explain the mobility mechanism of some active agents, like microbia. The way in which the active particle interacts with the boundary and the reorientation mechanism produce distinctive patterns of movement and distributions, giving us information about the underlaying laws governing its movement.
The introduction of process memory into the study of run-and-tumble particles reveals critical deviations from traditional theoretical predictions, emphasizing the need to consider historical dependencies in particle motion. This aligns with and expands upon findings in the broader literature, underscoring the complex interplay between memory, boundary interactions, and particle dynamics in confined settings.
This work not only affirms the robustness of the MPLT under a broad spectrum of conditions but also highlights the importance of angular distribution and boundary interactions in governing the dynamics of confined motion. The implications of our research extend beyond theoretical physics, offering insights into biological processes such as bacterial chemotaxis, and the design of microfluidic devices for sorting and analyzing microscopic particles.
8 Appendix
We present here an alternative procedure for calculating the radial distribution based on [46]. Consider the case of random angle boundary condition. The distribution can be calculated as follows. Consider a part of the trajectory inside a circle or radius $R$, as showed in Fig. (19). The particle is situated at the point $x$, at distance $r$ from the centre, and moves with direction $\overrightarrow{v}$, where $|\overrightarrow{v}|=1$. In this situation, the particle will follow the trajectory $\tau$, and will collide with the boundary at the point $y$. $v_{y}$ is the normal vector at the point $y$, as showed in Fig. (19).
Figure 19: Geometry of a single collision with the boundary.
Following [46], the probability of the particle of being in a given point with a certain direction is:
$\Pi(x,\theta)=\frac{\alpha}{|\overrightarrow{v}\cdot\overrightarrow{v_{y}}|},$ (25)
where $\alpha$ is a normalization constant. The scalar product of $\overrightarrow{v}$ and $\overrightarrow{v_{y}}$ is
$\overrightarrow{v}\cdot\overrightarrow{v_{y}}=|\overrightarrow{v}|.|\overrightarrow{v_{y}}|.cos\phi,$ (26)
and using the sine theorem,
$\frac{r}{sin\phi}=\frac{R}{sin\theta},$ (27)
we obtain
$|cos\phi|=\sqrt{1-sin^{2}\phi}=\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}.$ (28)
Therefore,
$\Pi(x,\theta)=\frac{\alpha}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}.$ (29)
The probability of being at the point $x$ is
$\rho(x)=\int_{0}^{2\pi}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=4\alpha\int_{0}^{\pi/2}\frac{\alpha d\theta}{\sqrt{1-\frac{r^{2}}{R^{2}}sin^{2}\theta}}=4\alpha K\left(\frac{r}{R}\right),$ (30)
where $K(x)$ is the complete elliptic integral of the first kind, defined as
$K(x)=\int_{0}^{\pi/2}\frac{ds}{\sqrt{1-x^{2}sin^{2}s}}.$ (31)
Finally, the radial probability is
$p(r)=2\pi r\cdot 4\alpha K\left(\frac{r}{R}\right).$ (32)
After normalization, we find that, $\alpha=\frac{1}{8\pi R^{2}}$, so
$p(r)=\frac{r}{R^{2}}K\left(\frac{r}{R}\right),$ (33)
the same results as in eq. (23). A similar calculation can be performed for the elastic case.
Acknowledgment
R.A. acknowledges an affiliation with INDAM.